# NRGPT: An Energy-based Alternative for GPT

### v1

@@ Abstract
11footnotetext: Equal contribution
{nima.dehmamy, benjamin.hoover, bishwajit.saha3, krotov}@ibm.com, leo_kozachkov@brown.edu, jjs@mit.edu
@@
Transformers represent a dominant paradigm in autoregressive language modeling (vaswani2017attention). In a typical setting, a sequence of tokens describing a text is passed through several transformer layers and mapped onto a new sequence, which is a copy of the original one shifted by one token and appended by the token that follows the initial sequence. At training time, this network is trained through self-supervised training, and at inference time the network is used for next token prediction. This is the standard Generative Pre-trained Transformer (GPT) setting, which is the first step in Large Language Model (LLM) design (radford2018improving).
@@
Energy-based modeling (lecun2006tutorial) is another prominent paradigm in modern AI landscape that historically goes back to Hopfield Networks (hopfield1982neural). In this framework the operation of the neural network is defined by a scalar energy function. Proper samples generated by the model (those that resemble training data) correspond to low energy states, while unrealistic samples (with large deviations from the training data distribution) correspond to high energy states. Thus EBMs are particularly appealing for their theoretical properties, because they view the “forward pass” through a deep network as an explicit optimization problem that maximizes the likelihood of the input data. Practically, this approach unlocks better tooling for systematic exploration of the solution space and enables natural solutions for both variable computation and model alignment via regularizers (see our extended discussion on EBMs for LLMs in Appendix˜C).
@@
Although at face value these two approaches look very different, in recent years a growing number of studies hint at deep connections. von2023transformers showed evidence that in-context learning (ICL) may be gradient descent by constructed an explicit weights such that the forward pass was GD on MSE loss. ahn2024transformers further showed that transformers learn a preconditioned GD for ICL. However, both of these works make significant simplifications, such as considering only linear transformers, omitting the softmax.
@@
Other works have have attempted to reconcile transformers and EBM from several angles. For instance, the Energy Transformer (hoover2023energy) is an architecture, which is simultaneously a transformer and an energy-based model. In the image domain, the typical setting would be to reconstruct a set of masked tokens (patches) given the set of open tokens. The network solves this task by performing a gradient descent of the energy on the space of tokens at inference time. This architecture is inspired by associative memory models (krotov2016dense) and for this reason solves the following problem: given a partially incomplete pattern – complete it in a meaningful way. This aspect of the core design makes it difficult to apply Energy Transformers to GPT settings, in which the sequence needs to be transformed to a shifted sequence by means of going through the network. Intuitively, in Energy Transformers the masked tokens need to evolve rapidly to match the missing parts of the pattern (e.g., image or graph), while the open tokens need to stay almost constant to barely adjust for the smooth transitions between the masked and the open tokens within the pattern. This is in drastic contrast with the GPT setting, in which there are no masked tokens at all. Rather, every token needs to evolve into the following token in the sequence.
@@
A different line of work is inspired by “System 2” thinking and attempts to design an energy-based network for processing language (gladstone2025energy). In this study, transformers are used as an architectural motif that casts text into a scalar energy function. While models of this nature have benefits for language processing, they belong exclusively to the class of energy-based models, and are unrelated to the GPT settings, commonly used in most LLMs.
@@
Individual modules within the transformer block, such as attention, have also been studied from the perspective of inference time optimization (geshkovski2023mathematical; geshkovski2024emergence). In this line of work, peculiar clustering properties of tokens have been observed. Energy-based optimization has also been studied in (yang2022transformers) from the perspective of majorization-minimization algorithms.
@@ 1 Energy-based modeling
In generative modeling our goal is to generate samples with a distribution close to observed datapoints. If we manage to learn an approximate likelihood function for the dataset, we can generate data by sampling. This is also the premise Energy-Based Models (EBM). An example of EBM would be Dense Associative Memory (krotov2016dense), where datapoints are stored in minima of an energy function. But more generally, the energy can represent a negative-log-likelihood, $E(\mathbf{x})=-\log P(\mathbf{x})$. In this case, the global minima of the energy represent maximum likelihood solutions. The deeper the energy, the higher the likelihood of that datapoint. One strategy to train an EBM is to first learn the energy function by fitting the distribution of the data. The sampling process would then be separate from learning the energy.
@@ 2 NRGPT Module
But subsequent works such as GPT-J (wang2021gpt), PaLM (chowdhery2023palm), and Energy Transformer (hoover2023energy) showed that the following parallel design has good performance too
@@ 2.1 Energy of NRGPT
$E$ $=E^{\mathrm{AT}}+E^{\mathrm{FF}},$ $\eta\partial_{g}E^{\mathrm{AT}}$ $=-\mathrm{AT}(\bm{g}),$ $\eta\partial_{g}E^{\mathrm{FF}}$ $=-\mathrm{FF}(\bm{g}),$ (7)
@@
Following hoover2023energy, define the attention energy
@@
In principle ${\bm{W}}^{V}$ and ${\bm{W}}^{P}$ can be merged into one matrix. (he2024simplifying) also experimented with removing ${\bm{W}}^{V}$ and ${\bm{W}}^{P}$ and found that in the setting without skip connections, these two weights could be largely omitted.
@@
The FF network generally has two layers $\mathrm{FF}(\bm{g}_{A})={{\bm{W}}^{2}}^{T}\sigma\left({\bm{W}}^{1}\bm{g}_{A}\right)$ with weights ${\bm{W}}^{1},{\bm{W}}^{2}\in\mathbb{R}^{M\times D}$, with $M$ being the size of the hidden layer. A possible choice for this network is a Dense Associative Memory (krotov2016dense). In this case
@@
NRGPT’s energy is distinct from prior works like that of Energy Based Transformers (EBT) (gladstone2025energy) and Energy Transformer (ET) (hoover2023energy). Specifically, EBT computes energies of next tokens an output of a standard transformer forward pass, while NRGPT describes a parameterized energy whose gradient returns a transformer block. Repeatedly minimizing the energy by following this gradient resembles a complete transformer forward pass. On the other hand, ET’s architecture appears more similar to NRGPT’s design, but it’s design was unsuited for causal token generation. We expand on these differences further in Appendix˜B.
@@ 2.2 Normalization of tokens
with $\bm{\mu}=\mathbb{E}[\bm{x}]$ for LayerNorm, and $\bm{\mu}=0,\bm{\delta}=0$ for RMSNorm. Here $\bm{\gamma},\bm{\delta}\in\mathbb{R}^{D}$ and $\odot$ is elementwise multiplication. Many recent models such as Qwen  bai2023qwen and Llama  grattafiori2024llama use RMSNorm  zhang2019root .
@@
Several works have shown careful initialization and scaling can replace normalization entirely (huang2020improving; wang2024deepnet). Doing so would remove much of the restrictions on the preconditioner $\eta$. In an $T$-layer transformer, Instead of layer normalization, T-Fixup (huang2020improving) and DeepNet (wang2024deepnet) initialize some weights and $x$ by a scale proportional to $T^{-1/4}$ and change most layer outputs to $f(g)\to\alpha f(x)$, with $\alpha\propto T^{-1/2}$. In this case, $\eta$ becomes much less restricted.
@@ 3.2 ListOps
We perform experiment on nested math operations on lists of integers, which are a version of ListOps (nangia2018listops). Our ListOps setting consists of three functions: maximum, median and sum modulo 20. Our inputs range from 0 to 19. Each data sample begins with nested equations like SUM(2,MAX(4,13,1),MEDIAN(5,3,16)). As performance metrics, we looked at accuracy on the mixed task, as well as the training and validation loss. Figure˜3 shows the results for two of our model variants, NRGPT_H_FF1 and NRGPT_H_FF2W as compared to GPT_Rec_parallel.
@@ 3.4 Open Web Text
Table˜2 shows the best model configuration for baseline GPT, RGPT-parallel, and our model NRGPT with the respective generation quality metrics. We see that the generation quality of NRGPT is very competitive with GPT and RGPT-parallel while it contains around 34M less parameters than GPT. Figure˜5 shows example of generated text by GPT, RGPT-parallel and NRGPT for which the generation quality metrics are provide in Table˜2. In these experiments, we consider transformer blocks configured at the same width — please see Section˜F.5 for more experiments across equal parameter counts, including experiments that show NRGPT’s advantage on downstream tasks such as MMLU (hendrycks2020measuring).
@@ 6 Reproducibility Statement
The code for our model, experiments, and analysis is available at https://github.com/bhoov/nrgpt in a self-contained environment. The code began as a fork of the excellent nanoGPT repositoryand as such all experiments and models are implemented using PyTorch (paszke2019pytorch). Each Shakespeare and ListOps experiment was conducted on a single H100 GPU. OWT experiments were conducted over 4 nodes of 8xH100 GPUs.
@@ B.1 Distinction from EBT
Both the Energy-Based Transformers (EBT) of (gladstone2025energy) and NRGPT are methods that can be used to minimize an explicit energy during autoregressive generation. However, the methods differ in how that energy is modeled. Specifically, the EBT paper uses standard, feed-forward transformer architectures with modified attention masks and a standard loss function on predicted tokens to turn the transformer predictions into an energy, whereas we use a novel, causal energy formulation of the transformer block inspired by the Energy Transformer (hoover2023energy).
@@ B.2 Distinction from ET
The architecture of Energy Transformer (ET) (hoover2023energy) is more similar to that of NRGPT, but it is fundamentally different. This is because ET is designed for MASKed-token prediction tasks whereas NRGPT is a special energy function designed for autoregressive token prediction, a paradigm that is not compatible with the original ET’s bidirectional attention design (where attention signal propagates both forward and backwards in time such that past tokens can attend to future tokens). A simple causal mask on ET attention breaks ET’s guarantee of monotonic energy minimization.
@@ Appendix C Advantages of EBMs
    Systematic exploration of the solution space. An explicit likelihood function enables us to systematically explore the space of solutions in LLMs using well-established methods in optimization and statistical physics, such as alternative gradient descent methods, minimum-energy paths, saddle-point analysis, and metastability. These tools are not available when the energy is implicitly encoded in a deep architecture, and we believe that finding alternative or creative solutions may correspond to exploring different local minima of the energy during inference. It also means that we can now view the forward process of causal LMs like GPT as a formal optimization process, which despite the prevalence of research around in-context learning (ICL) (dong2024survey), is not a widespread belief.
@@
    Variable computation using early stopping criteria. Both the energy and the norm of the energy’s gradient can be used as signals to stop model computation “early” for easier problems, or to continue thinking longer for harder ones. This is a primary motivation of other explicit EBM language models like EBT (gladstone2025energy), and we discuss this advantage further below.
@@
    Model alignment using energy regularizers. Note that Equation˜7 of our paper shows NRGPT’s architecture as the sum of two energies: a token-mixing attention energy $E^{\mathrm{AT}}$ and a token-wise feed-forward energy $E^{\mathrm{FF}}$. Per the precedent of ET (hoover2023energy), these are chosen to be faithful to the original transformer’s design. However, any scalar objective can theoretically be added to NRGPT’s block, and we imagine that adding regularizer terms to bias the energy landscape toward favorable solutions (or away from undesirable ones) is a unique benefit of EBMs that is more robust to prompt injection attacks than current LLMs.
@@ F.2 Evaluation Metrics
To assess the generation quality of our language models, we utilize several complementary metrics. We use Perplexity as a measure of model uncertainty, computed using a pretrained GPT-2 model to evaluate how well the generated text aligns with expected language patterns. Lower perplexity indicates more fluent and predictable text, with scores typically ranging from 10 (excellent) to 1000+ (poor quality). For lexical diversity, we utilize Distinct-1 and Distinct-2, which measure the ratio of unique unigrams and bigrams to total n-grams (or total words) in the generated text, respectively. These metrics range from 0 to 1, where higher values indicate greater vocabulary diversity and less repetitive generation. A Distinct-1 score near 0 suggests highly repetitive text, while scores above 0.8 indicate rich vocabulary usage. We utilize Average Pairwise Cosine Similarity using Sentence-BERT embeddings  (reimers2019sentence) to measure semantic diversity within generated samples. This metric calculates the mean cosine similarity between all pairs of generated sentences, ranging from -1 to 1. Optimal values fall between 0.3 and 0.6, balancing semantic diversity with topical coherence. Values below 0.3 indicate excessive divergence with potentially incoherent topic-jumping between sentences, while values above 0.7 suggest repetitive or redundant content with insufficient variation. The target range of 0.3-0.6 represents healthy diversity where generated sentences explore different aspects of a topic while maintaining semantic relevance and coherent narrative flow.
Finally, We compute the Grammar Quality Score (GQS), a composite metric that combines rule-based grammar error detection, spelling accuracy via spellchecker (pyspellchecker), and readability assessment using Flesch-Kincaid grade level (flesch1948new). GQS ranges from 0 (poor grammar) to 1 (perfect grammar), weighting grammatical correctness (50%), spelling accuracy (30%), and readability (20%). The metric identifies errors across ten categories including subject-verb agreement, tense consistency, and punctuation, with severity-weighted scoring. For complete context and to understand what the ideal ranges are for all of these metrics, see Table˜3.

### v2

@@ Abstract
Transformers represent a dominant paradigm in autoregressive language modeling (Vaswani et al., 2017). In a typical setting, a sequence of tokens describing a text is passed through several transformer layers and mapped onto a new sequence, which is a copy of the original one shifted by one token and appended by the token that follows the initial sequence. At training time, this network is trained through self-supervised training, and at inference time the network is used for next token prediction. This is the standard Generative Pre-trained Transformer (GPT) setting, which is the first step in Large Language Model (LLM) design (Radford et al., 2018).
@@
Energy-based modeling (LeCun et al., 2006) is another prominent paradigm in modern AI landscape that historically goes back to Hopfield Networks (Hopfield, 1982). In this framework the operation of the neural network is defined by a scalar energy function. Proper samples generated by the model (those that resemble training data) correspond to low energy states, while unrealistic samples (with large deviations from the training data distribution) correspond to high energy states. Thus EBMs are particularly appealing for their theoretical properties, because they view the “forward pass” through a deep network as an explicit optimization problem that maximizes the likelihood of the input data. Practically, this approach unlocks better tooling for systematic exploration of the solution space and enables natural solutions for both variable computation and model alignment via regularizers (see our extended discussion on EBMs for LLMs in Appendix˜C).
@@
Although at face value these two approaches look very different, in recent years a growing number of studies hint at deep connections. Von Oswald et al. (2023) showed evidence that in-context learning (ICL) may be gradient descent by constructed an explicit weights such that the forward pass was GD on MSE loss. Ahn et al. (2024) further showed that transformers learn a preconditioned GD for ICL. However, both of these works make significant simplifications, such as considering only linear transformers, omitting the softmax.
@@
Other works have have attempted to reconcile transformers and EBM from several angles. For instance, the Energy Transformer (Hoover et al., 2023) is an architecture, which is simultaneously a transformer and an energy-based model. In the image domain, the typical setting would be to reconstruct a set of masked tokens (patches) given the set of open tokens. The network solves this task by performing a gradient descent of the energy on the space of tokens at inference time. This architecture is inspired by associative memory models (Krotov and Hopfield, 2016) and for this reason solves the following problem: given a partially incomplete pattern – complete it in a meaningful way. This aspect of the core design makes it difficult to apply Energy Transformers to GPT settings, in which the sequence needs to be transformed to a shifted sequence by means of going through the network. Intuitively, in Energy Transformers the masked tokens need to evolve rapidly to match the missing parts of the pattern (e.g., image or graph), while the open tokens need to stay almost constant to barely adjust for the smooth transitions between the masked and the open tokens within the pattern. This is in drastic contrast with the GPT setting, in which there are no masked tokens at all. Rather, every token needs to evolve into the following token in the sequence.
@@
A different line of work is inspired by “System 2” thinking and attempts to design an energy-based network for processing language (Gladstone et al., 2025). In this study, transformers are used as an architectural motif that casts text into a scalar energy function. While models of this nature have benefits for language processing, they belong exclusively to the class of energy-based models, and are unrelated to the GPT settings, commonly used in most LLMs.
@@
Individual modules within the transformer block, such as attention, have also been studied from the perspective of inference time optimization (Geshkovski et al., 2023; 2024). In this line of work, peculiar clustering properties of tokens have been observed. Energy-based optimization has also been studied in (Yang et al., 2022) from the perspective of majorization-minimization algorithms.
@@ 1 Energy-based modeling
In generative modeling our goal is to generate samples with a distribution close to observed datapoints. If we manage to learn an approximate likelihood function for the dataset, we can generate data by sampling. This is also the premise Energy-Based Models (EBM). An example of EBM would be Dense Associative Memory (Krotov and Hopfield, 2016), where datapoints are stored in minima of an energy function. But more generally, the energy can represent a negative-log-likelihood, $E(\mathbf{x})=-\log P(\mathbf{x})$. In this case, the global minima of the energy represent maximum likelihood solutions. The deeper the energy, the higher the likelihood of that datapoint. One strategy to train an EBM is to first learn the energy function by fitting the distribution of the data. The sampling process would then be separate from learning the energy.
@@ 2 NRGPT Module
But subsequent works such as GPT-J (Wang and Komatsuzaki, 2021), PaLM (Chowdhery et al., 2023), and Energy Transformer (Hoover et al., 2023) showed that the following parallel design has good performance too
@@ 2.1 Energy of NRGPT
$E$ $=E^{\mathrm{AT}}+E^{\mathrm{FF}},$ $\bm{\eta}\partial_{g}E^{\mathrm{AT}}$ $=-\mathrm{AT}(\bm{g}),$ $\bm{\eta}\partial_{g}E^{\mathrm{FF}}$ $=-\mathrm{FF}(\bm{g}),$ (7)
@@
Following Hoover et al. (2023), define the attention energy
@@
In principle ${\bm{W}}^{V}$ and ${\bm{W}}^{P}$ can be merged into one matrix. (He and Hofmann, 2024) also experimented with removing ${\bm{W}}^{V}$ and ${\bm{W}}^{P}$ and found that in the setting without skip connections, these two weights could be largely omitted.
@@
The FF network generally has two layers $\mathrm{FF}(\bm{g}_{A})={{\bm{W}}^{2}}^{T}\sigma\left({\bm{W}}^{1}\bm{g}_{A}\right)$ with weights ${\bm{W}}^{1},{\bm{W}}^{2}\in\mathbb{R}^{M\times D}$, with $M$ being the size of the hidden layer. A possible choice for this network is a Dense Associative Memory (Krotov and Hopfield, 2016). In this case
@@
NRGPT’s energy is distinct from prior works like that of Energy Based Transformers (EBT) (Gladstone et al., 2025) and Energy Transformer (ET) (Hoover et al., 2023). Specifically, EBT computes energies of next tokens an output of a standard transformer forward pass, while NRGPT describes a parameterized energy whose gradient returns a transformer block. Repeatedly minimizing the energy by following this gradient resembles a complete transformer forward pass. On the other hand, ET’s architecture appears more similar to NRGPT’s design, but it’s design was unsuited for causal token generation. We expand on these differences further in Appendix˜B.
@@ 2.2 Normalization of tokens
with $\bm{\mu}=\mathbb{E}[\bm{x}]$ for LayerNorm, and $\bm{\mu}=0,\bm{\delta}=0$ for RMSNorm. Here $\bm{\gamma},\bm{\delta}\in\mathbb{R}^{D}$ and $\odot$ is elementwise multiplication. Many recent models such as Qwen  Bai et al. (2023) and Llama  Grattafiori et al. (2024) use RMSNorm  Zhang and Sennrich (2019) .
@@
Several works have shown careful initialization and scaling can replace normalization entirely (Huang et al., 2020; Wang et al., 2024). Doing so would remove much of the restrictions on the preconditioner $\eta$. In an $T$-layer transformer, Instead of layer normalization, T-Fixup (Huang et al., 2020) and DeepNet (Wang et al., 2024) initialize some weights and $x$ by a scale proportional to $T^{-1/4}$ and change most layer outputs to $f(g)\to\alpha f(x)$, with $\alpha\propto T^{-1/2}$. In this case, $\eta$ becomes much less restricted.
@@ 3.2 ListOps
We perform experiment on nested math operations on lists of integers, which are a version of ListOps (Nangia and Bowman, 2018). Our ListOps setting consists of three functions: maximum, median and sum modulo 20. Our inputs range from 0 to 19. Each data sample begins with nested equations like SUM(2,MAX(4,13,1),MEDIAN(5,3,16)). As performance metrics, we looked at accuracy on the mixed task, as well as the training and validation loss. Figure˜3 shows the results for two of our model variants, NRGPT_H_FF1 and NRGPT_H_FF2W as compared to GPT_Rec_parallel.
@@ 3.4 Open Web Text
Table˜2 shows the best model configuration for baseline GPT, RGPT-parallel, and our model NRGPT with the respective generation quality metrics. We see that the generation quality of NRGPT is very competitive with GPT and RGPT-parallel while it contains around 34M less parameters than GPT. Figure˜5 shows example of generated text by GPT, RGPT-parallel and NRGPT for which the generation quality metrics are provide in Table˜2. In these experiments, we consider transformer blocks configured at the same width — please see Section˜F.5 for more experiments across equal parameter counts, including experiments that show NRGPT’s advantage on downstream tasks such as MMLU (Hendrycks et al., 2020).
@@ 6 Reproducibility Statement
The code for our model, experiments, and analysis is available at https://github.com/bhoov/nrgpt in a self-contained environment. The code began as a fork of the excellent nanoGPT repository and as such all experiments and models are implemented using PyTorch (Paszke et al., 2019). Each Shakespeare and ListOps experiment was conducted on a single H100 GPU. OWT experiments were conducted over 4 nodes of 8xH100 GPUs.
Acknowledgments
The results presented here were obtained while Dmitry Krotov was employed by IBM Research. At the time of the camera-ready submission Dmitry Krotov is no longer employed by IBM Research.
@@ B.1 Distinction from EBT
Both the Energy-Based Transformers (EBT) of [Gladstone et al., 2025] and NRGPT are methods that can be used to minimize an explicit energy during autoregressive generation. However, the methods differ in how that energy is modeled. Specifically, the EBT paper uses standard, feed-forward transformer architectures with modified attention masks and a standard loss function on predicted tokens to turn the transformer predictions into an energy, whereas we use a novel, causal energy formulation of the transformer block inspired by the Energy Transformer [Hoover et al., 2023].
@@ B.2 Distinction from ET
The architecture of Energy Transformer (ET) [Hoover et al., 2023] is more similar to that of NRGPT, but it is fundamentally different. This is because ET is designed for MASKed-token prediction tasks whereas NRGPT is a special energy function designed for autoregressive token prediction, a paradigm that is not compatible with the original ET’s bidirectional attention design (where attention signal propagates both forward and backwards in time such that past tokens can attend to future tokens). A simple causal mask on ET attention breaks ET’s guarantee of monotonic energy minimization.
@@ Appendix C Advantages of EBMs
    Systematic exploration of the solution space. An explicit likelihood function enables us to systematically explore the space of solutions in LLMs using well-established methods in optimization and statistical physics, such as alternative gradient descent methods, minimum-energy paths, saddle-point analysis, and metastability. These tools are not available when the energy is implicitly encoded in a deep architecture, and we believe that finding alternative or creative solutions may correspond to exploring different local minima of the energy during inference. It also means that we can now view the forward process of causal LMs like GPT as a formal optimization process, which despite the prevalence of research around in-context learning (ICL) [Dong et al., 2024], is not a widespread belief.
@@
    Variable computation using early stopping criteria. Both the energy and the norm of the energy’s gradient can be used as signals to stop model computation “early” for easier problems, or to continue thinking longer for harder ones. This is a primary motivation of other explicit EBM language models like EBT [Gladstone et al., 2025], and we discuss this advantage further below.
@@
    Model alignment using energy regularizers. Note that Equation˜7 of our paper shows NRGPT’s architecture as the sum of two energies: a token-mixing attention energy $E^{\mathrm{AT}}$ and a token-wise feed-forward energy $E^{\mathrm{FF}}$. Per the precedent of ET [Hoover et al., 2023], these are chosen to be faithful to the original transformer’s design. However, any scalar objective can theoretically be added to NRGPT’s block, and we imagine that adding regularizer terms to bias the energy landscape toward favorable solutions (or away from undesirable ones) is a unique benefit of EBMs that is more robust to prompt injection attacks than current LLMs.
@@ F.2 Evaluation Metrics
To assess the generation quality of our language models, we utilize several complementary metrics. We use Perplexity as a measure of model uncertainty, computed using a pretrained GPT-2 model to evaluate how well the generated text aligns with expected language patterns. Lower perplexity indicates more fluent and predictable text, with scores typically ranging from 10 (excellent) to 1000+ (poor quality). For lexical diversity, we utilize Distinct-1 and Distinct-2, which measure the ratio of unique unigrams and bigrams to total n-grams (or total words) in the generated text, respectively. These metrics range from 0 to 1, where higher values indicate greater vocabulary diversity and less repetitive generation. A Distinct-1 score near 0 suggests highly repetitive text, while scores above 0.8 indicate rich vocabulary usage. We utilize Average Pairwise Cosine Similarity using Sentence-BERT embeddings  [Reimers and Gurevych, 2019] to measure semantic diversity within generated samples. This metric calculates the mean cosine similarity between all pairs of generated sentences, ranging from -1 to 1. Optimal values fall between 0.3 and 0.6, balancing semantic diversity with topical coherence. Values below 0.3 indicate excessive divergence with potentially incoherent topic-jumping between sentences, while values above 0.7 suggest repetitive or redundant content with insufficient variation. The target range of 0.3-0.6 represents healthy diversity where generated sentences explore different aspects of a topic while maintaining semantic relevance and coherent narrative flow.
Finally, We compute the Grammar Quality Score (GQS), a composite metric that combines rule-based grammar error detection, spelling accuracy via spellchecker [Barrus, 2020], and readability assessment using Flesch-Kincaid grade level [Flesch, 1948]. GQS ranges from 0 (poor grammar) to 1 (perfect grammar), weighting grammatical correctness (50%), spelling accuracy (30%), and readability (20%). The metric identifies errors across ten categories including subject-verb agreement, tense consistency, and punctuation, with severity-weighted scoring. For complete context and to understand what the ideal ranges are for all of these metrics, see Table˜3.